OpenStudy (anonymous):

Factor 7x^2 - 5x - 2

OpenStudy (anonymous):

I'm lost there's 2 primes :c

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

but this should help you solve it, i always hated factoring

OpenStudy (anonymous):

7x^2 - 5x - 2 = 7x^2 - 7x + 2x - 2 = 7x(x - 1) + 2(x - 1) = (7x + 2)(x - 1)